# In Defense of Defensive Forecasting

### v1

@@ 2 Rudiments of Defensive Forecasting for Predicting Events
Suppose we want to predict the likelihood that a certain event will occur based on observations of past events. For example, if we are going to predict whether a basketball player will make their next free throw, we will use their past success rate as a guess for the probability of the next shot. We can abstract this as observing a sequence of $T$ bits (assigning a $1$ if they make the shot and a 00 otherwise), and wanting to predict the likelihood that the next bit will be a $1$ (i.e., will they make the shot?). A reasonable heuristic guess for that likelihood is the average of the first $T$ bits. If there were far more ones than zeros, it is sensible to assume the next bit will also be more likely a one than not.
@@ 3 Defensive Forecasting: A Meta Algorithm
$\left\|\frac{1}{T}\sum_{t=1}^{T}F(x_{t},p_{t},y_{t})\right\|_{2}\leq\sqrt{\frac{\sum_{t=1}^{T}\left\|F(x_{t},p_{t},y_{t})\right\|_{2}^{2}}{T}}\leq\frac{M}{\sqrt{T}}\,.$
@@
In fact, all of the algorithms we derive here have the same form. We first find an efficiently computable function $S_{t}:[0,1]\rightarrow\mathbb{R}$ that summarizes the mistakes of the past. If $S_{t}(1)$ is nonnegative, we predict $1$. If not, we check $S_{t}(0)$. If it is non-positive, we predict 00 If neither condition holds, then we are guaranteed that $S_{t}(p)$ has a root in $[0,1]$. We find this root by binary search, and this root then serves as our prediction. We call this subroutine anticorrelation search, and summarize it in Algorithm 2.
@@ 4 Defensive Moment Matching
$\left\|\frac{1}{T}\sum_{i=t}^{T}(y_{t}-p_{t})\Phi(x_{t},p_{t})\right\|_{2}=\frac{C}{\sqrt{T}}\,.$ (6)
Then, for any function $f(x,p,y)$ such that, $f(x,p,1)-f(x,p,0)=\langle v,\Phi(x,p)\rangle$, we have
@@
Lemma 4.1 asserts that if Defensive Forecasting makes Equation 6 small, then we can effectively treat a large set of functions of $f(x,p,y)$ evaluated at the true outcomes as if the $y_{t}$ were sampled from a Bernoulli distribution with mean $p_{t}$. That is, for all intents and purposes, we can retrospectively pretend the $y_{t}$ are biased coin flips with the biases $p_{t}$ we wrote down as our predictions. Following Dwork et al. (2021) we will refer to predictions satisfying Equation 7 as being (online) outcome indistinguishable with respect to the set of functions $f$.
@@
where we used the assumption that discrete derivative $f(x,p,1)-f(x,p,0)$ is in the span of $\Phi$. We can bound this last term in terms of the norm of $v$ and a term made small by Defensive Forecasting.
@@
$\sup_{y\in\mathcal{Y}}\left\langle F(x_{t},p_{t},y),\sum_{s=1}^{t-1}F(x_{s},p_{s},y_{s})\right\rangle=\sup_{y\in\{0,1\}}(y-p_{t})\cdot S_{t}(p)\leq 0.$ (8)
@@ 5 Risk Minimization
$\sup_{a}\left|\sum_{t=1}^{T}\ell(a,y_{t})-\sum_{t=1}^{T}\mathbb{E}_{\tilde{y}\sim p_{t}}[\ell(a,\tilde{y}_{t})]\right|$ $\leq\mathcal{R}_{2}(T),$ (12)
@@
Note that this last result is a purely deterministic statement. It holds with probability 1 over the realized sequence of $y_{t}$. Furthermore, we made no assumptions (e.g. convexity) on the loss $\ell$ other than the fact that it is bounded and that $\ell(\pi(p_{t}),1)-\ell(\pi(p_{t}),0)$ is continuous in $p$.
@@ 6 Linear Classes and Online Learning
Now, let’s see how we can operationalize these ideas via Defensive Forecasting. To simplify notation and the algorithm, let’s focus our attention to the special case where $\ell$ is the squared loss, $\ell(p,y)=(y-p)^{2}$ and our action is simply predicting a $p_{t}\in[0,1]$. Let us assume that features $x$ are vectors in $\mathbb{R}^{d}$ with norm at most $B$, $\left\|x\right\|_{2}\leq B$, and that we are comparing against the class of linear predictors with norm at most $M$, =$\mathcal{H}=\{\langle x,w\rangle:\left\|w\right\|_{2}\leq M\}$.
@@
$\Phi(x,p)=\begin{bmatrix}1&p&x\end{bmatrix}^{\top}\,.$
@@
We note that there are other algorithms that achieve similar performance for linear prediction. Notably, the online gradient method that sets
$w_{t}$ $=w_{t-1}-\alpha(p_{t-1}-y_{t-1})x_{t}$
$p_{t}$ $=\langle w_{t},x_{t}\rangle$
will satisfy
@@
(see, for example, Theorem 1 in Zinkevich (2003)). As we can see by this expression, the incremental gradient method is also not making predictions about the future. Though we don’t know how to derive online gradient descent as a form of Defensive Forecasting, it isn’t too far away in spirit.
@@ 8 Calibration
[…]
@@
$\mathrm{Round}_{N}(p_{t})=\frac{1}{N}\begin{cases}\lfloor Np_{t}\rfloor\text{ with probability }\operatorname{frac}(Np_{t})\\ \lfloor Np_{t}\rfloor+1\text{ with probability }1-\operatorname{frac}(Np_{t})\\ \end{cases}$
where $\lfloor x\rfloor$ denotes the floor of $x$ and $\operatorname{frac}$ denotes the fractional part. Then with probability $1-\delta$ in the rounding procedure, we have for all integers $n$ between 00 and $N$
@@
[…]
Proof This operation is, in expectation, a smooth function of the prediction. In particular, for any integer $n$ between 00 and $N$,
@@
But we emphasize that beyond this property of shared interpretation, calibration doesn’t mean much. As many have emphasized before, a set of predictions can be perfectly calibrated and essentially useless for the purposes of decision-making. If a sequence of outcomes is a string of random bits with an equal number of $1$’s and 00’s, then predicting $p_{t}=1/2$ will be calibrated and it will be the best constant prediction as well. The value of any prediction algorithm is only good if its assumptions about the future turn out to be correct (say, that a constant prediction has low prediction error). Unfortunately, no algorithm can guarantee the future will look like the past.
@@ 10 Quantiles
Instead, for this particular evaluation metric, Defensive Forecasting is a deterministic algorithm that computes an approximation of the number $q$ by averaging 00s and $1$s. Let’s look at what happens in each round. Define,
@@
Then there are two cases. If $x_{t}\leq q$, we have by the inductive hypothesis $-\max\{q,1-q\}\leq e_{t}\leq 0$ and hence
@@ 11 Batch Learning with Defensive Forecasting
We can also extend online-to-batch conversion to our randomized Defensive Forecasting algorithm for predicting conditional quantiles. In this case, $\mathcal{A}_{t}(x_{t},\pi_{t})$ is a randomized procedure. Hence, we can consider $f_{t}(x_{t})$ to be a distribution $\Delta_{t}^{p}$ over $[Y_{\mathrm{min}},Y_{\mathrm{max}}]$ that is supported on 2 points. Still, all the analysis thus far works, we just define our batch prediction for $x$ by first sampling an $f_{i}$ uniformly and then sampling $p$ from $f_{i}(x)$. When $f_{i}$ are distributions over predictions, we write,
@@
Statisticians often refer to these results as conditional coverage statements since the probability that $y$ is greater than $p$ is equal to $1-q+o(1)$ conditional on the event that $f(x,p)=1$.
@@ Acknowledgements
The authors would like to thank Siva Balakrishnan, Unai Fischer-Abaigar, Dean Foster, Laurent Lessard, Tengyuan Liang, Aaron Roth, and Ryan Tibshirani for several helpful discussions, comments, and suggestions. JCP was generously supported in part by the Harvard Center for Research on Computation and Society and the Alfred P. Sloan Foundation Grant G-2020-13941. BR was generously supported in part by NSF CIF award 2326498, NSF IIS Award 2331881, and ONR Award N00014-24-1-2531.

### v2

@@ 2 Rudiments of Defensive Forecasting for Predicting Events
Suppose we want to predict the likelihood that a certain event will occur based on observations of past events. For example, if we are going to predict whether a basketball player will make their next free throw, we will use their past success rate as a guess for the probability of the next shot. We can abstract this as observing a sequence of $T$ bits (assigning a $1$ if they make the shot and a 0 otherwise), and wanting to predict the likelihood that the next bit will be a $1$ (i.e., will they make the shot?). A reasonable heuristic guess for that likelihood is the average of the first $T$ bits. If there were far more ones than zeros, it is sensible to assume the next bit will also be more likely a one than not.
@@ 3 Defensive Forecasting: A Meta Algorithm
$\left\|\frac{1}{T}\sum_{t=1}^{T}F(x_{t},p_{t},y_{t})\right\|_{2}\leq\sqrt{\frac{\sum_{t=1}^{T}\left\|F(x_{t},p_{t},y_{t})\right\|_{2}^{2}}{T^{2}}}\leq\frac{M}{\sqrt{T}}\,.$
@@
In fact, all of the algorithms we derive here have the same form. We first find an efficiently computable function $S_{t}:[0,1]\rightarrow\mathbb{R}$ that summarizes the mistakes of the past. If $S_{t}(1)$ is nonnegative, we predict $1$. If not, we check $S_{t}(0)$. If it is non-positive, we predict 0 If neither condition holds, then we are guaranteed that $S_{t}(p)$ has a root in $[0,1]$. We find this root by binary search, and this root then serves as our prediction. We call this subroutine anticorrelation search, and summarize it in Algorithm 2.
@@ 4 Defensive Moment Matching
$\left\|\frac{1}{T}\sum_{i=t}^{T}(y_{t}-p_{t})\Phi(x_{t},p_{t})\right\|_{2}\leq\frac{C}{\sqrt{T}}\,.$ (6)
Then, for any function $f(x,p,y)$ such that, $f(x,p,1)-f(x,p,0)=\langle v,\Phi(x,p)\rangle$ where $v$ is an arbitrary fixed vector, we have
@@
Lemma 4.1 asserts that if Defensive Forecasting makes Equation 6 small, then we can effectively treat a large set of functions of $f(x,p,y)$ evaluated at the true outcomes as if the $y_{t}$ were sampled from a Bernoulli distribution with mean $p_{t}$. That is, for all intents and purposes, we can retrospectively pretend the $y_{t}$ are biased coin flips with the biases $p_{t}$ we wrote down as our predictions. Following Dwork et al. (2021) we will refer to predictions satisfying Equation 7 as being (online) outcome indistinguishable with respect to the set of functions $f$. Later on in sections Section 5 and Section 8 we will describe various interesting classes of functions $f$ that can be written this way, $f(x,p,1)-f(x,p,0)=\langle v,\Phi(x,p)\rangle$.
@@
where we used the assumption $f(x,p,1)-f(x,p,0)=\langle v,\Phi(x,p)\rangle$. We can bound this last term in terms of the norm of $v$ and a term made small by Defensive Forecasting.
@@
$\sup_{y\in\mathcal{Y}}\left\langle F(x_{t},p_{t},y),\sum_{s=1}^{t-1}F(x_{s},p_{s},y_{s})\right\rangle=\sup_{y\in\{0,1\}}(y-p_{t})\cdot S_{t}(p_{t})\leq 0.$ (8)
@@ 5 Risk Minimization
$\left|\sum_{t=1}^{T}\ell(\pi(p_{t}),y_{t})-\sum_{t=1}^{T}\mathbb{E}_{\tilde{y}_{t}\sim p_{t}}[\ell(\pi(p_{t}),\tilde{y}_{t})]\right|$ $\leq\mathcal{R}_{1}(T)$ (11)
$\sup_{a}\left|\sum_{t=1}^{T}\ell(a,y_{t})-\sum_{t=1}^{T}\mathbb{E}_{\tilde{y}_{t}\sim p_{t}}[\ell(a,\tilde{y}_{t})]\right|$ $\leq\mathcal{R}_{2}(T),$ (12)
@@
Note that this last result is a purely deterministic statement. It holds with probability 1 over the realized sequence of $y_{t}$. Furthermore, we made no assumptions (e.g. convexity) on the loss $\ell$ other than the fact that it is bounded and that $\ell(\pi(p_{t}),1)-\ell(\pi(p_{t}),0)$ is continuous in $p$. Furthermore, for simplicity, we considered the case where there are no context vectors $x_{t}$. However, the same ideas generalize to that setting as we will see in a moment.
@@ 6 Linear Classes and Online Learning
Now, let’s see how we can operationalize these ideas via Defensive Forecasting. To simplify notation and the algorithm, let’s focus our attention to the special case where $\ell$ is the squared loss, $\ell(p,y)=(y-p)^{2}$ and our action is simply predicting a $p\in[0,1]$. Let us assume that features $x$ are vectors in $\mathbb{R}^{d}$ with norm at most $B$, $\left\|x\right\|_{2}\leq B$, and that we are comparing against the class of linear predictors with norm at most $M$, $\mathcal{H}=\{\langle x,w\rangle:\left\|w\right\|_{2}\leq M\}$.
@@
$\Phi(x,p)=\begin{bmatrix}1&p&x\end{bmatrix}^{\top}\,$. (15)
@@
We note that there are other algorithms that achieve similar performance for linear prediction. Notably, given a step size parameter $\alpha>0$, the online gradient method sets,
$w_{t}$ $=w_{t-1}-\alpha(p_{t-1}-y_{t-1})x_{t-1},$
and predicts $p_{t}=\langle w_{t},x_{t}\rangle$. This sequence of predictions achieves,
@@
Setting $\alpha$ appropriately, the online gradient method has a similar $\sqrt{T}$ excess risk bound. See, for example, Theorem 1 in Zinkevich (2003). As we can see by this expression, the incremental gradient method is also not making predictions about the future. Though we don’t know how to derive online gradient descent as a form of Defensive Forecasting, it isn’t too far away in spirit or in functional form.
@@ 8 Calibration
Moreover, for any $\alpha\in[0,1]$ and $\epsilon\leq\min(\alpha,1-\alpha)$, define $h_{\epsilon,\alpha}(p)=w_{\epsilon}(p-\alpha)$. The functions $h_{\epsilon,\alpha}$ map $[0,1]$ to $[0,1]$ and have small Fermi-Sobolev norm because $\int_{-1}^{1}w_{\epsilon}(t)dt=\epsilon$ and $\int_{-1}^{1}|w_{\epsilon}^{\prime}(t)|^{2}dt=2/\epsilon$, so we have
@@
[…]
where $\lfloor x\rfloor$ denotes the floor of $x$ and $\operatorname{frac}$ denotes the fractional part. Then with probability $1-\delta$ in the rounding procedure, we have for all integers $n$ between 0 and NN
@@
Note that the rounding algorithm here maps the prediction to its closest point on a grid with spacing $1/N$. For instance, if $N=20$, the rounding operation maps $p_{t}=.89$ to $v^{\prime}=.9$ with probability .2.2 and to $v=.85$ with probability .8.8. If $N$ is chosen to be equal to $T^{1/3}$, the corresponding bound has a regret of $T^{2/3}$ that is the regret accrued by most sequential calibration algorithms (Dagan et al., 2025).
Proof This operation is, in expectation, a smooth function of the prediction. In particular, for any integer $n$ between 0 and $N$,
@@
But we emphasize that beyond this property of shared interpretation, calibration doesn’t mean much. As many have emphasized before, a set of predictions can be perfectly calibrated and essentially useless for the purposes of decision-making. If a sequence of outcomes is a string of random bits with an equal number of $1$’s and 0’s, then predicting $p_{t}=1/2$ will be calibrated and it will be the best constant prediction as well. The value of any prediction algorithm is only good if its assumptions about the future turn out to be correct (say, that a constant prediction has low prediction error). Unfortunately, no algorithm can guarantee the future will look like the past.
@@ 10 Quantiles
Instead, for this particular evaluation metric, Defensive Forecasting is a deterministic algorithm that computes an approximation of the number $q$ by averaging 0s and $1$s. Let’s look at what happens in each round. Define,
@@
Then there are two cases. If $x_{t}\leq q$, we have by the inductive hypothesis $\max\{q,1-q\}\leq e_{t}\leq 0$ and hence
@@ 11 Batch Learning with Defensive Forecasting
We can also extend online-to-batch conversion to our randomized Defensive Forecasting algorithm for predicting conditional quantiles. In this case, $\mathcal{A}_{t}(x_{t},\pi_{t})$ is a randomized procedure. Hence, we can consider $f_{t}(x_{t})$ to be a distribution $\Delta_{t}^{p}$ over $Y_{\mathrm{min}},Y_{\mathrm{max}}]$ that is supported on 2 points. Still, all the analysis thus far works, we just define our batch prediction for $x$ by first sampling an $f_{i}$ uniformly and then sampling $p$ from $f_{i}(x)$. When $f_{i}$ are distributions over predictions, we write,
@@
Statisticians often refer to these results as conditional coverage statements since the probability that $y$ is greater than $p$ is equal to 1q+o(1)1-q+o(1) conditional on the event that $f(x,p)=1$.
@@ Acknowledgements
The authors would like to thank Siva Balakrishnan, Unai Fischer-Abaigar, Dean Foster, Laurent Lessard, Tengyuan Liang, Christopher Musco, Aaron Roth, and Ryan Tibshirani for several helpful discussions, comments, and suggestions. JCP was generously supported in part by the Harvard Center for Research on Computation and Society and the Alfred P. Sloan Foundation Grant G-2020-13941. BR was generously supported in part by NSF CIF award 2326498, NSF IIS Award 2331881, and ONR Award N00014-24-1-2531.